sir, in trigo phase 3 can you tell the proof for:-
acotA+bcotB+ccotC=2(R+r)

Grade:11

1 Answers

s arvind
9 years ago
from sine rule 
sinA=a/2R
sinB=b/2R
sinc=c/2R    
   
substituting the value in the above equation,we get
 
2R(cosA +cosB +cosC)
 
using conditional identities,
 
cosA + cosB + cosC = 1+4sinA/2 sinB/2 sinC/2 = 1+ r/R
 
hence,2R(1+r/R) is otained